OpenStudy (anonymous):

A man 180cm tall casts a shadown 150cm long. The shadow of a flagpole is 1050cm long. How tall is the flagpole?

OpenStudy (bahrom7893):

Use similar triangles

OpenStudy (bahrom7893):

180cm and 150cm xcm and 1050cm

OpenStudy (bahrom7893):

180/150=x/1050 x = 180*1050/150

OpenStudy (bahrom7893):

x=1260